OpenStudy (anonymous):

Help if possible!! What are the approximate solutions of 2x2 + 9x = 8 to the nearest hundredth? (1 point) x ≈ 1.22 and x ≈ 3.28 x ≈ -0.76 and x ≈ 5.26 x ≈ -1.22 and x ≈ -3.28 x ≈ -5.26 and x ≈ 0.76

OpenStudy (anonymous):

option 2

OpenStudy (anonymous):

now i believe that you start with adding on both sides and then dividing

OpenStudy (anonymous):

oops option 4

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

Use the quadratic formula to get the answer.

OpenStudy (anonymous):

and that other question poojah and i worked on for 30 mins was wrong

OpenStudy (anonymous):

\[x = \frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\] \[2x ^{2}+9x -8 =0\] a = 2 b = 9 c =-8 Plug in the numbers and try it out

OpenStudy (anonymous):

\[2x^2 + 9x = 8\\2x^2 + 9x-8=0\\standard\ quadratic\ eq.\ ax^2+bx+c=0\\ on\ comparing\, we\ have\ a=2, b=9, c=-8\\using\ quadratic\ formula\\x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\\ put\ the \ values\ of\ a,\ b\ c\\we\ have\ x=\frac{ -9 \pm \sqrt{9^2-4.2.(-8)} }{ 2.2 }\\x=\frac{ -p \pm \sqrt{81+64} }{ 4 }\\x= \frac{ -9 \pm \sqrt{145} }{ 4}\\x=\frac{ -9 \pm 12.04 }{ 4 }\\x=\frac{ -9-12.04 }{ 4},\ \frac{ -9+12.04 }{ 4}\\x=\frac{ -21.04 }{ 4},\ \frac{ 3.04 }{ 4}\\x=-5.26,\ x=0.76[ nearest\ hundredth]\]

OpenStudy (anonymous):

\[please,\ replace\ = by\ \approx\ as \sqrt{145} \ has \ been\ taken\ approximately=12.04\\(nearesst\ hundredth)\]